How many products can be formed from the numbers 2, 3, 4, 5, 6, 7 by taking two or more numbers at a
time?
1
Expert's answer
2020-03-04T10:02:43-0500

"n=\{2,3,4,5,6,7\}"

"k_1=2;\;k_2=3."

Amount of products of the numbers is given by


"C_n^{k_1}+C_n^{k_2}=\frac{n!}{k_1!(n-k_1)!}+\frac{n!}{k_2!(n-k_2)!}""=\frac{6!}{2!4!}+\frac{6!}{3!3!}=35."